OpenStudy (anonymous):

Solve the equation 5x-1/9 = 9/5x-1?

OpenStudy (precal):

is that (1/9) or is that (5x-1)/9?

OpenStudy (anonymous):

\[\frac{ 5x-1 }{ 9 }=\frac{ 9 }{ 5x-1 }\]

OpenStudy (anonymous):

yeap

OpenStudy (precal):

ok cross multiply

OpenStudy (anonymous):

What he said.

OpenStudy (precal):

what she said

OpenStudy (precal):

(5x-1)(5x-1)=9(9) expand out and finish solving

OpenStudy (precal):

you might want to factor or use the quadratic formula to solve for x